Active mass of 5.6L of N2 at STP
1
Expert's answer
2020-05-01T14:21:58-0400

According to the definition, the active mass of nitrogen is:

"c = \frac{n}{V}" ,

where "n" is the number of the moles of nitrogen and "V" is the volume of the system. It is the same as molar concentration. Let's find the number of the moles of nitrogen in 5.6 L at STP (pressure "p" =105 Pa, temperature "T"= 273.15 K) . Assuming that nitrogen has the behavior of the ideal gas:

"pV = nRT"

"n = \frac{pV}{RT}" ,

where "R" is the ideal gas constant, equal to 8.314·103 L Pa / (K mol).

"n = \frac{10^5·5.6}{8.314 ·10^3·273.15} = 0.247" mol

Finally, the active mass of nitrogen is:

"c = \frac{0.247}{5.6} = 0.044" mol/L

Answer: Active mass of 5.6L of N2 at STP is 0.044 mol/L